The alcohol content of hard liquor is normally given in terms of the “proof,” which is defined as twice the percentage by volume of ethanol (C2H5OH) present. Calculate the concentration (M) of alcohol that is in 1.20 L of 60-proof gin. The density of ethanol is 0.798 g/mL


1
Expert's answer
2021-05-28T01:29:55-0400

We have given the density of alcohol "= 0.798"

We know,

"Density = \dfrac{Mass}{Volume}"


"0.798 = \dfrac{M}{1200}"


"M = 957.6g"

Now we can calculate the number of moles


"n = \dfrac{M}{m}"


"Molarity = \dfrac{n}{V}"


Hence, concentration can be calculated.